How many grams of solid potassium cyanide should be added to 1.00 L of a 0.279 M hydrocyanic acid solution to prepare a buffer with a pH of 10.100? The Ka of hydrocyanic acid is 4.00×10-10.

Mass of potassium cyanide =
1
Expert's answer
2020-11-02T08:46:02-0500

Pka = -log(ka) = -log(4.00×10-10) = 9.4

for buffer solution , henderson equation is PH = Pka + log([salt]\[acid])

so or 10.1 = 9.4 +log ([KCN]\0.279)

or log([KCN]\0.279) = 10.1 - 9.4 = 0.7

or [KCN]\0.279 = 100.7 = 5.0

or [KCN] = 5.0 x 0.279 = 1.395

so molarity or concentration of KCN = 1.395 M

so we have to add 1.395 mol of KCN in 1 L solution of HCN.

molar mass of KCN = 65.12 g mol-1

so mass of KCN required = 1.395 x 65.12 = 90.84 g of KCN. ( ans)